solve each of the following systems by graphing
x-2y=8
3x-2y=12

Answer by venugopalramana(3286)

SEE
\solve each of the follwing systems by substitution.
5x-2y=-5............I
y-5x=3......................II...Y=3+5X..........PUT IN EQN.I
5X-2(3+5X)=-5
5X-6-10X=-5
-5X=6-5=1
X=-1/5
Y=3+5X=3+5*(-1/5)=3-1=2
THESE ARE CONSISTENT INDEPENDENT EQNS.AND HENCE HAVE A UNIQUE SOLUTUIN NAMELY...
=-1/5 AND Y=2THE FOLLOWING AND TRY